The theoretical yield of a reaction is the amount of product obtained if the limiting reactant is completely converted to product.

Consider the reaction:
CH4(g) + CCl4(g) → 2 CH2Cl2(g)

If 17.38 g CH4 is mixed with 13.50 g CCl4, calculate the theoretical yield (g) of CH2Cl2 produced by the reaction.
Answer #1

Molar mass of CH4,

MM = 1*MM(C) + 4*MM(H)

= 1*12.01 + 4*1.008

= 16.042 g/mol

mass(CH4)= 17.38 g

use:

number of mol of CH4,

n = mass of CH4/molar mass of CH4

=(17.38 g)/(16.04 g/mol)

= 1.083 mol

Molar mass of CCl4,

MM = 1*MM(C) + 4*MM(Cl)

= 1*12.01 + 4*35.45

= 153.81 g/mol

mass(CCl4)= 13.5 g

use:

number of mol of CCl4,

n = mass of CCl4/molar mass of CCl4

=(13.5 g)/(1.538*10^2 g/mol)

= 8.777*10^-2 mol

Balanced chemical equation is:

CH4 + CCl4 ---> 2 CH2Cl2

1 mol of CH4 reacts with 1 mol of CCl4

for 1.083 mol of CH4, 1.083 mol of CCl4 is required

But we have 8.777*10^-2 mol of CCl4

so, CCl4 is limiting reagent

we will use CCl4 in further calculation

Molar mass of CH2Cl2,

MM = 1*MM(C) + 2*MM(H) + 2*MM(Cl)

= 1*12.01 + 2*1.008 + 2*35.45

= 84.926 g/mol

According to balanced equation

mol of CH2Cl2 formed = (2/1)* moles of CCl4

= (2/1)*8.777*10^-2

= 0.1755 mol

use:

mass of CH2Cl2 = number of mol * molar mass

= 0.1755*84.93

= 14.91 g

Answer: 14.91 g